### v1

# Decoding Noise in Nanofluidic Systems: Adsorption versus Diffusion Signatures in Power Spectra

Abstract
Adsorption processes play a fundamental role in molecular transport through nanofluidic systems, but their signatures in measured signals are often hard to distinguish from other processes like diffusion. In this paper, we derive an expression for the power spectral density (PSD) of particle number fluctuations in a channel, accounting for diffusion and adsorption/desorption to a wall. Our model, validated by Brownian dynamics simulations, is set in a minimal yet adaptable geometry, allowing us to eliminate the effects of specific geometries. We identify distinct signatures in the PSD as a function of frequency $f$, including $1/f^{3/2}$ and $1/f^{1/2}$ scalings related to diffusive entrance and re-entrance effects, and a $1/f^{2}$ scaling associated with adsorption. These scalings appear in key predicted quantities – the total number of particles in the channel and the number of adsorbed or unadsorbed particles – and can dominate or combine in non-trivial ways depending on parameter values. Notably, when there is a separation of timescales between diffusion within the channel and adsorption/desorption times, the PSD can exhibit two distinct slopes in some of the predicted quantities. We provide a phase diagram to classify experimental systems based on predicted PSD shapes. These PSDs reflect measured properties in physical systems on the nano- and micro-scale, such as ion channels, nanopores, and electrochemical sensors, potentially offering insights into noisy experimental data.
preprint: AIP/123-QED
I Introduction
The diffusive passage of particles in confined geometries is a key phenomenon in many different systems – see Fig. 1. A classic example is transport through protein channels, which facilitates the flow of ions and larger biomolecules across the cell membrane Neher and Sakmann (1976); Cooper, Jakobsson, and Wolynes (1985); Nikaido (1992). Inspired by this, protein channels have been used as a biotechnological tool for nanopore-based sequencing devices Kasianowicz et al. (1996); Clarke et al. (2009); Bell and Keyser (2016); Deamer, Akeson, and Branton (2016). Moreover, both biological and artificial nanopores have been increasingly developed for applications ranging from water filtration and blue energy harvesting Werber, Osuji, and Elimelech (2016); Siria, Bocquet, and Bocquet (2017) to functionalized single-molecule or single-particle sensing devices Ying et al. (2022); Dekker (2007); Goto et al. (2020); Siwy, Bruening, and Howorka (2023); Chazot-Franguiadakis et al. (2022). A common feature of the measured signals in these systems is a relatively high level of noise. This arises, in part, because only small numbers of particles are involved at such small scales, but also arises due to the diffusive nature of particle motion Kavokine, Netz, and Bocquet (2021); Verveen and DeFelice (1974).
Figure 1: Adsorption and diffusion phenomena in nano- and micro- scale systems. In the schematics, particles (black) diffuse under confinement, with their relevant lengthscales highlighted. In these systems attractive interactions like adsorption (red) between the particles and the confining walls are also important, and add a degree of complexity.
In many of the previous cases, attractive interactions, like adsorption between the transported particles and the confining surface itself play a key role (see red interfaces in Fig. 1). For example, protein channels make use of specific binding to encode selectivity Luckey and Nikaido (1980); Benz, Schmid, and Vos-Scheperkeuter (1987); Andersen, Jordy, and Benz (1995), and in nanopore sensing technologies, adsorption has been used to prolong the dwell time of species in the channel, which aids detection Howorka, Cheley, and Bayley (2001); Wei, Tampé, and Rant (2012); Yusko et al. (2011); Rotem et al. (2012); Wiggin et al. (2008); Jetha et al. (2011). Adsorption can also transiently modify redox reactions for electrochemical detection Singh et al. (2011). In these cases, adsorption processes often act as an additional source of experimental noise. Adsorption to the pore walls may cause transient blockage of a nanopore and thus fluctuations in the measured ionic current Andersen, Jordy, and Benz (1995); Nekolla, Andersen, and Benz (1994). It may also lead to fluctuations in the channel’s surface charge, and hence in its ionic conductivity Hoogerheide, Garaj, and Golovchenko (2009); Secchi et al. (2016).
While detrimental from a sensing perspective, attempts have been made to use the link between characteristic noise signatures and molecular processes to reveal physical properties of the underlying system. For example, in nanopores, noise signatures could help identify the number of ionizable sites in a pore Bezrukov and Kasianowicz (1993) or the presence of analyte or cation adsorption to the walls Knowles et al. (2021); Singh et al. (2011); Queralt-Martín, Perini, and Alcaraz (2021); Gravelle, Netz, and Bocquet (2019). In nanoscale electrochemical devices, similar techniques have determined how reversible analyte adsorption to electrodes affects current noise Kätelhön et al. (2014); Zevenbergen et al. (2009).
A methodology for understanding noisy signals common to many experimental systems is to investigate the power spectral density (PSD) of the measured signal. These investigations have revealed $1/f^{\alpha}$ scalings, where $\alpha\simeq 0.5-2$ and $f$ is the frequency, in biological and synthetic nanopores Smeets et al. (2008); Smeets, Dekker, and Dekker (2009); Knowles, Keyser, and Thorneywork (2019); Fuliński et al. (1998); Fragasso, Pud, and Dekker (2019); Powell et al. (2009); Wohnsland and Benz (1997); Siwy and Fuliński (2002) and electrochemical sensing systems Zevenbergen et al. (2009); Kätelhön et al. (2014). Many theoretical approaches have revealed similar scaling laws, allowing us in principle to connect theories and experiments Bezrukov et al. (2000); Berezhkovskii, Pustovoit, and Bezrukov (2002); Marbach (2021); Gravelle, Netz, and Bocquet (2019); Zevenbergen et al. (2009); Nestorovich et al. (2002); Zorkot, Golestanian, and Bonthuis (2016a, b). Direct comparison is, however, limited to simple and specific cases. For instance, real experimental systems often exhibit heterogeneous transport properties at the single particle level Knowles, Mackay, and Thorneywork (2024), blurring the PSD signals. Often several noisy processes operate simultaneously, and models are used which require several fitting parameters, whose physical interpretation is limited Robin et al. (2023a). Thus, in many cases theoretical approaches are still needed to help make concrete links between such scaling laws and molecular mechanism.
In particular, it is difficult to rationalize the coupled effects of adsorption and diffusive transport. Previous theoretical work has investigated the effect on the PSD of a single binding site in a cylindrical channel Berezhkovskii, Pustovoit, and Bezrukov (2002) and adsorption to the walls of a pore in combination with the effect of reservoirs Gravelle, Netz, and Bocquet (2019); Robin et al. (2023a). In these models, it can be difficult to distinguish the effects of adsorption from other model-specific features affecting particle dynamics, especially the geometry, which introduces several time- and length-scales Marbach (2021); Robin et al. (2023a). This has led some authors to linearly sum PSDs induced by uniquely diffusion or adsorption to account for experimental data Zevenbergen et al. (2009), a likely flawed attempt in systems where the timescales for both phenomena can be similar Knowles et al. (2021); Gadgil et al. (2004). In particular, in experimental data involving adsorption and diffusion, it is still unclear when adsorption yields specific signatures in the PSD, as it is sometimes apparently linked to the emergence of two distinct slopes Zevenbergen et al. (2009); Robin et al. (2023a) and sometimes not Knowles et al. (2021); Secchi et al. (2016).
To answer this question, we introduce a simple framework to explore and separate the effects of diffusion and adsorption on power spectra. Our model involves non-interacting particles diffusing along a one-dimensional line, with a segment where particles can adsorb and become immobile – see Fig. 2. This model is versatile and can easily be adapted to various geometries, such as channels Marbach (2021) or electrochemical devices Zevenbergen et al. (2009). Its simplicity allows us to isolate geometric effects from those of diffusion and adsorption. We derive analytical expressions, validated by simulations, for the auto-correlation function and PSD of particle number fluctuations for all particles in the channel, as well as for free (unadsorbed) and bound (adsorbed) particles (Sec. II). The resulting PSDs are complex combinations of the purely diffusive and adsorbing processes, showing distinct scalings like $1/f^{1/2}$, $1/f^{3/2}$, and $1/f^{2}$, which combine non-linearly. We rationalize the mechanisms behind each of these scaling laws: diffusive re-entrance and entrance effects and adsorption/desorption phenomena, respectively. Crucially, when there is significant difference in timescales between diffusion across the channel and adsorption/desorption, the PSDs of some but not all of the predicted quantities exhibit two distinct slopes (Sec. III). Finally, we discuss strategies to solve the inverse problem of determining adsorption properties from experimental data (Sec. IV).
II Theoretical Model
II.1 Setup to Probe the Effect of Adsorption
We consider a 1D system of non-interacting particles diffusing with diffusion coefficient $D$ along a line $P$ with length $L$ – see Fig. 2. In the center of the line, a region of length $L_{0}$ represents a “channel region”. Inside the channel, in addition to diffusing, the particles can adsorb to a separate line $Q$ with rate $k_{\rm on}$. On line $Q$, the particles cannot diffuse, but remain “stuck”. Particles can unbind from $Q$ to line $P$ with rate $k_{\rm off}$. Binding and unbinding rates are independent of the presence of other particles on either line. In this equilibrium scenario, the ratio of the rates corresponds to a Boltzmann factor, as $k_{\rm on}/k_{\rm off}\simeq e^{-e_{0}/k_{B}T}$ where $e_{0}$ is the effective adsorption energy and $k_{B}T$ the unit of energy. When adsorption is favorable, $e_{0}\leq 0$, and $k_{\rm on}\geq k_{\rm off}$. The probability density of being in position $x$ at time $t$ on line $P$ is given by $p(x,t)$, and on line $Q$ by $q(x,t)$.
Figure 2: A representation of our model system of a channel connected to two reservoirs (line $P$) and adsorption on the channel wall (line $Q$). The channel has length $L_{0}$ while the total length of the system is $L$. Adsorption and desorption rates are set by $k_{\rm on}$ and $k_{\rm off}$, respectively.
In choosing such a general 1D system rather than a specific geometry, our model is easily adaptable to different geometries. For example, it can be applied to systems with two binding regions on either side of the domain, as in many electrochemical cells, or a single binding region on one side, as in test assays. We can also change $L$ and $L_{0}$ at will to resemble systems where the binding region is large ($L\sim L_{0}$, as in electrochemical devices and assays) or small ($L\gg L_{0}$, as in many nanopore systems). Some of these different geometries are illustrated in Fig. 1.
Our goal is to establish the statistical properties of the number of particles inside the channel, $N(t)$, and how this number is split between free ($N_{\rm off}(t)$) and adsorbed ($N_{\rm on}(t)$) particles. From the fluctuations of these numbers, we will calculate correlation functions and their associated power spectra. Each of these quantities – freely diffusing particles, adsorbed particles, and their total – may be of interest in different experimental settings, as different experimental methods may be sensitive to different particle subsets.
To make progress on analytics, we write the system of equations obeyed by $p(x,t)$ and $q(x,t)$ as:
$\begin{cases}&\mathrm{for}\,\,|x|\leq L_{0}/2,\,\displaystyle\begin{cases}&\partial_{t}p=-k_{\rm on}p+k_{\rm off}q+D\partial_{xx}p\\ &\partial_{t}q=+k_{\rm on}p-k_{\rm off}q\end{cases}\\ &\mathrm{for}\,\,|x|>L_{0}/2,\,\partial_{t}p=D\partial_{xx}p.\end{cases}$ (1)
The boundary conditions are (i) continuity in $p$ at the edges of the channel, i.e. at $x=\pm L_{0}/2$, and (ii) no flux into the walls at the edges of the domain, i.e. $\partial_{x}p=0$ at $x=\pm L/2$. Note that the relevant continuous variable at the edges of the channel is $p$, rather than $p+q$, because there is no mechanism that allows particles to go directly from the reservoirs to being adsorbed inside the channel in one infinitesimal time step.
To quantify statistical properties, we need to establish the probabilities of transitioning from one state to another. For this reason, we start by considering a situation where a particle is initially located in the channel. It can begin either adsorbed, on line $Q$, or unadsorbed, on line $P$. Noting that the partition function of the system is $Z=L+L_{0}(k_{\rm on}/k_{\rm off})$, we can calculate the initial probability densities
$p_{0}$ $=p(|x|\leq L_{0}/2,t=0)=\frac{1}{Z}$ (2)
$q_{0}$ $=q(|x|\leq L_{0}/2,t=0)=\frac{k_{\rm on}/k_{\rm off}}{Z}$
for these two cases, and $p(|x|\geq L_{0}/2,t=0)=0$.
The system of equations in Eq. (1) can now be solved, working in Laplace space; full details of the calculation are in Appendix A. From this solution we obtain expressions for $\hat{p}(x,s)$ and $\hat{q}(x,s)$, the Laplace transforms of variables $p(x,t)$ and $q(x,t)$, as
$\hat{p}=A_{C}e^{m_{1}x}+B_{C}e^{-m_{1}x}+\frac{k_{\rm off}(p_{0}+q_{0})+sp_{0}}{s(s+k_{\rm off}+k_{\rm on})}$ (3)
and
$\hat{q}=\frac{k_{\rm on}}{s+k_{\rm off}}(A_{C}e^{m_{1}x}$ $+B_{C}e^{-m_{1}x})$ (4)
$+$ $\frac{k_{\rm on}(p_{0}+q_{0})+sq_{0}}{s(s+k_{\rm off}+k_{\rm on})}$
where $s$ is the Laplace frequency, $m_{0}$ and $m_{1}$ are functions of $s$, and $A_{C}=B_{C}$ is a constant determined from the boundary conditions. Note that the expression for $\hat{p}$ is only true inside the channel where $|x|\leq L_{0}/2$.
In Sec. II.2, we use these results to evaluate the total number fluctuations in the channel and their correlation function. Following this, in Sec. II.3 we determine the power spectral density of the fluctuations.
II.2 Number Fluctuations in the Channel
In the following section, we calculate the fluctuating number of particles in the channel, $N(t)$, and then evaluate the corresponding correlation function. The same method can be used to calculate the correlation functions for the number of adsorbed particles $N_{\rm on}(t)$ and the number of unadsorbed particles $N_{\rm off}(t)$; details are in Appendix B.
We express the correlation function as
$C_{N}(t)=\langle N(t)N(0)\rangle=\mathcal{N}\psi_{C\rightarrow C}(t)$ (5)
where $\psi_{C\rightarrow C}(t)$ represents the probability that a particle starting in the channel is still in the channel at time $t$, and $\mathcal{N}$ is the total number of particles in the system. Note that the correct prefactor is indeed $\mathcal{N}$, the total number of particles in the system rather than the mean number in the channel. Formally, this originates from the fact that we are doing a canonical ensemble average here. Physically, any particle in the system has a probability to be in the channel at time $t=0$ and so we need to find the correlation function most generally for all of them – and since particles are non-interacting, we can simply multiply the probability by the total number of particles inside the system.
Since we took initial conditions such that all particles started in the channel, we find $\psi_{C\rightarrow C}(t)$ simply by integrating the probability densities $p(x,t)$ and $q(x,t)$ over the channel length
[…]
We take the time derivative to obtain an expression for the correlation function as
[…]
$=2\mathcal{N}\int_{0}^{L_{0}/2}D\partial_{xx}p(x,t)dx$
$=2\mathcal{N}D[\partial_{x}p(x,t)]^{L_{0}/2}_{0}$
and taking the Laplace transform we arrive at
$\hat{C}_{N}(s)=2\mathcal{N}\frac{m_{1}}{m_{0}}A_{C}(e^{L_{0}m_{1}/2}-e^{-L_{0}m_{1}/2}).$ (8)
II.3 Spectral Density
The power spectrum $S_{N}(f)$ of number fluctuations at a given frequency $f$ is an important tool in analyzing their behavior at different timescales and is commonly used to characterize noisy experimental data in nanofluidic systems. It can be calculated from the correlation function of the number of particles in the channel using the Wiener-Khinchin theorem Bendat and Piersol (2010)
[…]
$S_{N}(f)=4\Re[\hat{C_{N}}(2i\pi f)],$ (10)
with $\Re[\cdot]$ the real part of a scalar and where the additional factor of 2 comes from the fact that the power spectral density is calculated for positive frequencies only, $S_{>}(\omega)=2S(\omega)$, so that the total power is conserved Bendat and Piersol (2010). Therefore, using the result in Eq. (8) above, we find for the power spectrum of the number fluctuations in the channel
$S_{N}(f)=8\Re\left[\mathcal{N}\frac{m_{1}}{m_{0}^{2}}A_{C}\Big{(}e^{L_{0}m_{1}/2}-e^{-L_{0}m_{1}/2}\Big{)}\right].$ (11)
Substituting $A_{C}$, and applying the same transformation to obtain the power spectrum of the number of bound or free particles in the channel, we obtain
$\begin{gathered}S_{N}(f)=8\Re\Bigg{[}\frac{\mathcal{N}m_{1}(k_{\rm off}(p_{0}+q_{0})+sp_{0})}{sm_{0}(s+k_{\rm off}+k_{\rm on})}\frac{(e^{Lm_{0}}-e^{L_{0}m_{0}})(1-e^{L_{0}m_{1}})}{(m_{0}-m_{1})(e^{Lm_{0}}-e^{L_{0}(m_{0}+m_{1})})+(m_{0}+m_{1})(e^{Lm_{0}+L_{0}m_{1}}-e^{L_{0}m_{0}})}\Bigg{]},\\ S_{\rm on}(f)=8\Re\Bigg{[}\frac{\mathcal{N}k_{\rm off}q_{0}}{s(s+k_{\rm off}+k_{\rm on})}\Bigg{(}-\frac{L_{0}}{2}+\frac{m_{0}k_{\rm on}}{m_{1}(s+k_{\rm off})}\frac{(e^{Lm_{0}}-e^{L_{0}m_{0}})(1-e^{L_{0}m_{1}})}{(m_{0}-m_{1})(e^{Lm_{0}}-e^{L_{0}(m_{0}+m_{1})})+(m_{0}+m_{1})(e^{Lm_{0}+L_{0}m_{1}}-e^{L_{0}m_{0}})}\Bigg{)}\Bigg{]},\\ S_{\rm off}(f)=8\Re\Bigg{[}\frac{\mathcal{N}k_{\rm on}p_{0}}{s(s+k_{\rm off}+k_{\rm on})}\Bigg{(}-\frac{L_{0}}{2}+\frac{m_{0}(s+k_{\rm off})}{m_{1}k_{\rm on}}\frac{(e^{Lm_{0}}-e^{L_{0}m_{0}})(1-e^{L_{0}m_{1}})}{(m_{0}-m_{1})(e^{Lm_{0}}-e^{L_{0}(m_{0}+m_{1})})+(m_{0}+m_{1})(e^{Lm_{0}+L_{0}m_{1}}-e^{L_{0}m_{0}})}\Bigg{)}\Bigg{]}.\end{gathered}$ (12)
Note that $s=2i\pi f$ in the above expression. Eq. (12) is the main theoretical result of this paper. In its current form, it is hardly amenable to physical interpretation, however, and our goal in the following sections is to uncover limiting regimes and mechanistic behaviour from this expression.
III Results
In spite of its apparent simplicity, our system is governed by four key timescales:
  • Time to diffuse across the system: $\tau_{\rm sys}=L^{2}/2D$
  • Time to diffuse across the channel: $\tau_{\rm cha}=L_{0}^{2}/2D$
  • Adsorption timescale: $\tau_{\rm ads}=1/k_{\rm on}$
  • Desorption timescale: $\tau_{\rm des}=1/k_{\rm off}$
These timescales can be estimated for many nanotechnological systems and we report orders of magnitude in Table 1. Remarkably, there is no standard case, and relative magnitudes can vary strongly from one system to another.
Example $\tau_{\rm cha}$ (s) $\tau_{\rm ads}$ (s) $\tau_{\rm des}$ (s)
Case 1: $\tau_{\rm cha}\gg\tau_{\rm ads}\sim\tau_{\rm des}$
Antibody/antigen binding in a microfluidic chip Hu, Gao, and Li (2007) $10^{5}$ $10^{2}$ $10^{3}$
Molecules reacting in nanofluidic cavities Zevenbergen et al. (2009) $10^{0}$ $10^{-1}$ $10^{-3}$
Case 2: $\tau_{\rm cha}\ll\tau_{\rm ads}\sim\tau_{\rm des}$
Sugar transport in transmembrane proteins Nekolla, Andersen, and Benz (1994); Kullman, Winterhalter, and Bezrukov (2002) $10^{-8}$ $10^{-2}$ $10^{-3}$
Adapter adsorption inside nanopores Gu, Cheley, and Bayley (2001) $10^{-5}$ $10^{-2}$ $10^{-3}$
Nano-electrochemical devices Kätelhön et al. (2014) $10^{-7}$ $10^{-4}$ $10^{-5}$
Case 3: $\tau_{\rm cha}\sim\tau_{\rm ads}\sim\tau_{\rm des}$
PEG adsorption in glass nanopores Knowles et al. (2021) $10^{-4}$ $10^{-4}$ $10^{-2}$
DNA hybridization in microarrays Gadgil et al. (2004) $10^{3}$ $10^{2}$ $10^{4}$
Case 4: $\tau_{\rm ads}\gg\tau_{\rm cha}\sim\tau_{\rm des}$
Biomolecule sensing in porous silicon membranes Zhao et al. (2016) $10^{0}$ $10^{-1}$ $10^{6}$
Gas phase catalysis in zeolites Hansen et al. (2009) $10^{-2}$ $10^{-2}$ $10^{2}$
Table 1: Typical values for the timescales of channel diffusion, adsorption, and desorption, in a number of different experimental systems. We do not include $\tau_{\rm sys}$ here; it is harder to determine in general and in most cases will either be similar to the channel length ($L\sim L_{0}$) or much greater ($L\gg L_{0}$, in which case the limit of an infinite system is approached).
In the following, we start by investigating limit cases of pure diffusion or only adsorption. This will allow us to establish a few canonical fluctuation regimes before turning to the full, much more complex system.
Related to this, we work in the following reduced units. In all cases the characteristic length scale is $\ell_{0}=L_{0}$, and the characteristic timescale is $\tau_{0}=\tau_{\rm cha}=L_{0}^{2}/2D$, except in the case of pure adsorption/desorption where $D=0$, in which case we use $\tau_{0}=1/f_{0}$, where $f_{0}=(k_{\rm on}+k_{\rm off})/2\pi$ and represents the corner frequency in the PSD. We compare the analytical results, derived above, to particle-based simulations of the system (see Appendix C for details on computational methods).
III.1 Pure Diffusion
We first verify that our results recover the expected behavior for the case of pure diffusion of particles in a channel Bezrukov et al. (2000); Marbach (2021). This scenario corresponds to setting $k_{\rm on}=0$ and $q_{0}=0$. Here, we obtain $S_{\rm on}(f)=0$, $S_{\rm off}(f)=S_{N}(f)$ and
$S_{N}(f)=4\Re\Bigg{[}\frac{\mathcal{N}p_{0}}{m_{0}s}\frac{(e^{Lm_{0}}-e^{L_{0}m_{0}})(e^{-L_{0}m_{0}}-1)}{(e^{Lm_{0}}-1)}\Bigg{]}$ (13)
where we recall that in the expression above $s=2i\pi f$.
The theoretical (line) and simulation (dots) spectra for the total number of particles are shown in Fig. 3 (purple) and are in perfect agreement. The spectrum clearly involves multiple regimes; more specifically, a plateau at low frequencies and then two scaling regimes with exponents $-1/2$ and $-3/2$. The scaling regimes can be interrogated by considering the limits of Eq. (13).
[…]
Firstly, as $f\rightarrow 0$ we find:
$S_{N}(f)$ $=\frac{\mathcal{N}p_{0}(L-L_{0})^{2}L_{0}^{2}}{3DL}$ (14)
i.e. there is a plateau at low frequencies. At high frequencies as $f\rightarrow\infty$, we find instead that the PSD tends to
$S_{N}(f)$ $=\frac{\mathcal{N}p_{0}\sqrt{D}}{(\pi)^{3/2}}\frac{1}{f^{3/2}}$ (15)
with a scaling as $1/f^{3/2}$. Between these two regimes, the slope of the PSD appears to vary steadily between the two limits; however, this masks a well-defined underlying scaling. To find this, we take $L\rightarrow\infty$ and $f\rightarrow 0$
$S_{N}(f)$ $=\frac{\mathcal{N}p_{0}L_{0}^{2}}{\sqrt{D\pi}}\frac{1}{\sqrt{f}}$ (16)
predicting a scaling as $1/f^{1/2}$. Physically, we can interpret these different regimes as follows.
Plateau is linked to the mean correlation time.
For low frequencies, corresponding to timescales longer than the time for a particle to diffuse across the system $\tau_{\rm sys}$, we see a plateau (white noise). This is due to the finite size of the system – fluctuations in particle number are eventually uncorrelated because the particles have had sufficient time to equilibrate outside of the channel by diffusion. The plateau value of a PSD spectrum corresponds to a characteristic timescale as $S_{N}(f)=\mathcal{N}\tau_{\rm corr}$ where $\tau_{\rm corr}$ is the mean correlation time of the number fluctuations,
$\tau_{\rm corr}=\int_{0}^{\infty}\frac{C_{N}(t)}{\mathcal{N}}\mathrm{d}t.$ (17)
Remembering that $p_{0}=1/L$ and taking $L\gg L_{0}$ for simplicity, we find $\tau_{\rm corr}=\frac{L_{0}^{2}}{3D}=2\tau_{\rm cha}/3$. The plateau amplitude thus depends on the mean time particles spend inside the channel. It is interesting that the amplitude does not depend on system size $L$ but actually probes the channel size $L_{0}$, and the only dependence of the plateau on $L$ is the frequency at which it begins.
High frequency $1/f^{3/2}$ is linked to diffusive particle crossings in and out of the channel.
At high frequencies, corresponding to timescales shorter than the time for a particle to diffuse across the channel, $\tau_{\rm cha}$, we see a characteristic $1/f^{3/2}$ scaling. This has been observed in previous works Marbach (2021); Lax and Mengert (1960); Zevenbergen et al. (2009); Berezhkovskii, Pustovoit, and Bezrukov (2002); Minh, Rotenberg, and Marbach (2023). Its origin is particles crossing from one region to another (i.e. from channel to reservoir and vice versa), by diffusion. This result is independent of the system geometry.
Intermediate $1/f^{1/2}$ is linked to re-entrance effects.
The intermediate regime between these two timescales – where we observe a slope in the PSD lower than $3/2$, but have not yet reached a plateau – is a signature of the underlying $1/f^{1/2}$ scaling. While its presence has been highlighted in Ref. [Marbach, 2021], its origin has not been identified. Here, we show this scaling arises from re-entrance effects, i.e. particles diffusing back into the channel that have previously diffused out of the channel at intermediate timescales. This leads to long-lived correlations that produce an additional scaling regime in the PSD before correlations are removed over timescales longer than $\tau_{\rm sys}$.
As further evidence of this, we use simulations to probe a modified version of our model system as defined in Fig. 2. Here, the particles diffuse normally along the line $P$, up to the point that they cross the channel boundaries. When a particle leaves the channel, however, it is re-allocated to a location randomly picked in the reservoir. In this way, we artificially remove the “return” correlation in particle trajectories. Indeed, this entirely removes the scaling regime as $1/f^{1/2}$ between the channel and system timescales, as shown in Fig. 3 (green dots). As a result, the onset of the plateau regime occurs around a frequency corresponding to the time to diffuse across the channel $1/\tau_{\rm cha}$, and not at $1/\tau_{\rm sys}$; by removing the correlations when particles exit the channel, they are no longer probing the finite system size, but only the finite channel size.
III.2 Only Adsorption/Desorption
The second key source of fluctuations in the full problem is the adsorption/desorption process itself. To consider characteristics of this in isolation, i.e. adsorption in a system with no diffusion, we set $D=0$ in the original system of equations Eq. (1) and solve the system for $p$ and $q$. Here, there will naturally be no fluctuations in the total number of particles in the channel region ($S_{N}(f)=0$), but there are still fluctuations in the number of bound and free particles. As both quantities are directly related to one another in that case, their associated spectra are equal, $S_{\rm on}(f)=S_{\rm off}(f)$. Additionally, we can use $p_{0}$ and $q_{0}$ interchangeably as they are related by $p_{0}k_{\rm on}=q_{0}k_{\rm off}$. In this situation, we obtain
$S_{\rm on/off}(f)=\frac{\mathcal{N}_{\rm off}k_{\rm on}}{\pi^{2}}\frac{1}{f^{2}+f_{0}^{2}}$ (18)
where $\mathcal{N}_{\rm off}=\mathcal{N}L_{0}p_{0}$ is the mean number of free particles, and we recall that $f_{0}=2\pi/(k_{\rm off}+k_{\rm on})$. This expression is consistent with works which investigate exclusively adsorption/desorption processes Zevenbergen et al. (2009); Nestorovich et al. (2002).
Spectra obtained from Eq. (18) and simulations for the pure adsorption/desorption scenario are compared in Fig. 4.
Figure 4: Power spectral density of number fluctuations with no diffusion and only adsorption/desorption processes. Spectrum for fluctuations in the number of unadsorbed particles is shown, which is equivalent to that for adsorbed particles. Dashed lines indicate $\tau_{\rm ads}=1/k_{\rm on}$ and $\tau_{\rm des}=1/k_{\rm off}$. A solid line indicates $f_{0}$, the corner frequency, which is the mean of the two rates $f_{0}=(k_{\rm on}+k_{\rm off})/2$. There is a clear $1/f^{2}$ scaling at high frequencies and a plateau below $f_{0}$. In this case, $L_{0}=10\ell_{0}$, $k_{\rm on}=0.009f_{0}$, and $k_{\rm off}=2f_{0}$. We use separated adsorption and desorption rates to demonstrate the dependence of the corner primarily on the faster rate (here, $k_{\rm off}$).
The shape of the spectrum is a Lorentzian, scaling as $1/f^{2}$ at high frequencies. This can be intuitively understood as, at very short timescales, the fluctuating signal appears as individual step functions. Transforming into frequency space, a step function becomes $1/f$, which is then squared to obtain the PSD and $1/f^{2}$ scaling. At low frequencies, or long timescales, the fluctuating signal is “smeared out” and resembles a white noise, or a plateau in the PSD. Remarkably, the binding/unbinding process corresponds to a single shot noise process, which again intuitively makes sense because binding and unbinding events are uncorrelated and thus follow a Poisson distribution.
The cross-over frequency between the two regimes is $f_{0}=(k_{\rm on}+k_{\rm off})/2\pi$, i.e. the mean of the adsorption and desorption rates. This can be understood by considering the situation where one of the two processes becomes very fast. For e.g. fast desorption $k_{\rm off}\gg k_{\rm on}$, the probability of an adsorption event being followed later in time by another adsorption event is independent of the lifetime of the adsorbed state, becoming a pure shot noise process in the limit of zero adsorption time. As such, correlations in the noise only persist for frequencies larger than the faster rate. In contrast, as lifetimes on and off the surface become comparable, the probability of adsorption will be influenced by previous adsorption events, resulting in a corner frequency that is the average of the two rates. The plateau value again corresponds to a mean correlation time, as $S_{\rm on/off}(f\rightarrow 0)=\mathcal{N}_{\rm off}q_{0}\tau_{0}$, where the relevant correlation time is $\tau_{0}=1/(2\pi f_{0})=1/(k_{\rm on}+k_{\rm off})$, the mixture between the two timescales.
Similar noise processes and therefore spectral signatures are observed in various other physical systems. For example, this noise is similar to random telegraph or “popcorn” noise – noise observed in semiconductors where there are sudden transitions between two (or sometimes more) current levels Machlup (1954). In semiconductors, this is often due to the generation and trapping of charge carriers. Analogously, here, this noise is due to particles having only two states, “on” or adsorbed, and “off” or unadsorbed. Equally, the process can be mapped to scenarios where noise arises from signals involving “free particle pulses”. Here, the “pulse” is the mean number of free particles $\mathcal{N}_{\rm off}$ and occurs with mean frequency $k_{\rm on}$, which corresponds to the mean time particles stay free. Considering that $\mathcal{N}_{\rm off}k_{\rm on}=\mathcal{N}_{\rm on}k_{\rm off}$, the process can also be interpreted as “bound particle pulses”. The subtlety is that the amplitude of each of these pulses decays exponentially with a timescale which is the mixture of the two timescales at play $\tau_{0}$. The exponential distribution of the pulses is a specificity of this system, and thus, compared to other processes whose length in time are constant, as in Ref. [Knowles, Mackay, and Thorneywork, 2024], the shape of the PSD has no oscillations and is a true Lorentzian Verveen and DeFelice (1974).
Figure 5: Slow diffusion, fast adsorption/desorption: Case 1 – $\tau_{\rm cha}\gg\tau_{\rm ads}\sim\tau_{\rm des}$. Left: total PSD. Middle: bound PSD. Right: free PSD. We use $L=2\ell_{0}$, $k_{\rm on}=20/\tau_{0}$, and $k_{\rm off}=60/\tau_{0}$. Note here that $k_{\rm on}\leq k_{\rm off}$. The dotted curve indicates the pure diffusive case with the same parameters (but $k_{\rm on}=0$). Similarly, the dash-dotted curve indicates the pure adsorption/desorption case with the same parameters (but $D=0$). Relevant timescales are labeled and marked with dashed vertical lines.
III.3 Full Case
We now consider the full problem where both diffusion and adsorption processes lead to fluctuations. Notably, in this case the results from the previous two sections do not add linearly. Instead, we observe a much more complicated picture, both in the analytical expressions in Eq. (12) and in the shapes and behavior of the PSDs, although some distinguishable “common” scaling regimes are reached. To build intuition, we therefore consider the following limiting cases:
  1. 1.
    Slow diffusion, fast adsorption/desorption: $\tau_{\rm sys}\sim\tau_{\rm cha}\gg\tau_{\rm ads}\sim\tau_{\rm des}$
  2. 2.
    Fast diffusion, slow adsorption/desorption: $\tau_{\rm sys}\sim\tau_{\rm cha}\ll\tau_{\rm ads}\sim\tau_{\rm des}$
  3. 3.
    All processes have comparable timescales: $\tau_{\rm sys}\sim\tau_{\rm cha}\sim\tau_{\rm ads}\sim\tau_{\rm des}$
  4. 4.
    Fast adsorption compared to desorption: $\tau_{\rm des}\gg\tau_{\rm sys}\sim\tau_{\rm cha}\gg\tau_{\rm ads}$
We refer to the power spectra of the total particles in the channel, the adsorbed particles, and the free particles in the channel as the “total PSD”, “bound PSD”, and “free PSD”, respectively. The first two cases we consider in detail in the main text, and the second two in Appendix D. Note that in the following, the dimensionless timescale $\tau_{0}=L_{0}^{2}/2D$ refers to the time to diffuse across the channel.
III.3.1 Diffusion Slower than Adsorption/Desorption
We first explore the case where diffusion is slower than both adsorption and desorption processes. Physically, slow diffusive timescales (relative to reaction kinetics) often appear in microfluidic systems. For example, the kinetics of antigen/antibody binding is fast compared to the time for antigens to diffuse across an immunoassay Hu, Gao, and Li (2007). Similarly, the adsorption/desorption of redox-active species in nanofluidic electronic devices can be fast with respect to their diffusion Zevenbergen et al. (2009).
Results (both theory and simulations) are shown in Fig. 5, where the corresponding curves for a pure diffusion (dotted black line) or adsorption (dash-dotted black line) process are shown for comparison. The total PSD is close to the purely diffusive case, with a corner frequency around the diffusion timescale and a high frequency scaling of $-3/2$. It is perhaps counter-intuitive that the total PSD, $S_{N}$, does not show the corner frequency linked to adsorption given that the pure adsorption curve has a higher magnitude than the diffusion curve at high frequencies. This is a result of cross-correlations, which may be negative, i.e., even though $N=N_{\rm on}+N_{\rm off}$, the power spectra are not additive, $S_{N}\neq S_{\rm on}+S_{\rm off}$.
Bound and free PSDs are similar to each other and both display two corner frequencies, one around the diffusive timescales and one at the binding timescales. With respect to scalings, the bound PSD appears to follow the pure adsorption case at high frequencies, while the free PSD deviates from this and seems to have a slightly less steep slope than $-2$ for very high frequencies, possibly even approaching the $-3/2$ regime. At low frequencies, both free and bound PSDs plateau at values that are below that of the purely diffusive case, again highlighting non-additivity. This is likely a scaling effect associated with the average number of bound/free particles being less than that of the particles in the channel. In the intermediate regime between the two corners, there is a transient scaling regime; this appears as $-3/2$ due to diffusion for the free PSD, but is sufficiently short that this may not always be obvious in experimental systems.
The bound PSD is very similar to the spectrum of fluctuations in the current measured in Ref. [Zevenbergen et al., 2009]. In that nanofluidic system, it is likely the relevant quantity being probed by the current is the number of species adsorbed onto the electrodes, which means our results for the bound PSD are in good agreement. Our model rationalizes these results in the context of other fluctuating quantities in the system. Finally, we note that the ordering of adsorption and desorption timescales does not significantly affect the shape of the power spectra; the complementary case where adsorption is faster than desorption ($k_{\rm on}\gg k_{\rm off}$) is discussed in Appendix D, in particular in Fig. 8.
III.3.2 Diffusion Faster than Adsorption/Desorption
We now turn to the situation where diffusion is faster than adsorption/desorption. Physically, this ordering of timescales is common in nanopores, both biological and engineered. For example, it occurs naturally in the transport of sugars through maltoporins (LamB transmembrane proteins) Kullman, Winterhalter, and Bezrukov (2002); Nekolla, Andersen, and Benz (1994) and also in the use of adapters, such as beta cyclodextrin, in nanopore sensors for detection of organic molecules Gu, Cheley, and Bayley (2001). Nanoscale electrical devices where analytes reversibly bind to the electrodes may also involve fast diffusion relative to the adsorption timescales Kätelhön et al. (2014).
Representative spectra are shown in Fig. 6. Here, we only consider the case where $k_{\rm on}=k_{\rm off}$, because varying $k_{\rm on}$ and $k_{\rm off}$ does not significantly affect the shape or behavior of the PSDs, provided both correspond to timescales much shorter than that for diffusion.
Figure 6: Fast diffusion, slow adsorption/desorption: Case 2 – $\tau_{\rm cha}\ll\tau_{\rm ads}\sim\tau_{\rm des}$ Left: total PSD. Middle: bound PSD. Right: free PSD. We use $L=2\ell_{0}$, $k_{\rm on}=k_{\rm off}=0.008/\tau_{0}$. The color/line code is the same as that of Fig. 5.
We observe that the total PSD is more sensitive to the adsorption/desorption process compared to the previous case. In particular, it now shows two corners and two slopes, with the low-frequency corner determined by the adsorption/desorption timescale and high frequency corner by the diffusive timescale(s). Again, there is a transient intermediate scaling regime between the two corners. The slopes of all three PSDs at high frequencies seem to take the same values as in the pure adsorption or pure diffusion cases respectively (i.e. $-2$ and $-3/2$), depending on the PSD.
In this case, the total and free PSDs have similar shapes whereas previously the free and bound PSDs were similar. Here, with fast diffusion, the number of bound particles is less dependent on the diffusive timescale as a rate-limiting step. Instead, at high frequencies and short timescales, the number of bound particles remains roughly constant, and the total number of particles reflects changes in the number of free particles. This results in a $-3/2$ scaling in total and free PSDs due to particles diffusing in and out of the reservoirs. At low frequencies, the free and total PSDs show corners at the adsorption/desorption timescale, since at longer times particles can equilibrate across the whole system (diffusing, adsorbing, and unadsorbing) and so the fluctuations become uncorrelated.
In the context of Ref. [Zevenbergen et al., 2009], their observation of $-3/2$ scaling at intermediate frequencies and $-2$ at high frequencies does not appear in this case in any of the measured fluctuating quantities. As such, we can conclude that the diffusive timescales must be slower than the adsorption/desorption timescales for their system. This highlights that the ordering of slopes in the PSD can be indicative of the relative timescales in the system, a relatively simple criterion to observe.
III.3.3 Further Cases
Figure 7: Left: physical systems corresponding to different diffusion/adsorption regimes. Right: the corresponding shapes of the PSDs identified in this work.
As highlighted in Table 1, other experimental systems may exhibit further different timescale separations, where all timescales are comparable (Case 3) or where adsorption is much faster than all other processes (Case 4). In these scenarios, the spectra obtained are again non-linear combinations of the pure diffusion and pure adsorption cases, and can either be quite similar to either of these pure cases (Case 3) or exhibit entirely peculiar shapes (Case 4). More detailed investigations are reported in Appendix D.
To summarize the different possibilities for the PSD shapes, we build a phase diagram – see Fig. 7, right-hand side – comparing the diffusive and the adsorption/desorption timescales respectively. A key takeaway is that there exists only a few special cases in which the PSDs exhibit two corners, related to two distinct slopes, while in all other cases, the PSDs appear to have an approximately Lorentzian, smooth, shape, with a single well-defined or somewhat smeared out scaling regime. These special “two distinct slopes” cases are only observed in case 1 of slow diffusion compared to binding timescales, for bound and free PSDs; and in case 2 of fast diffusion compared to binding timescales, for total and free PSDs. The free PSD thus always senses the two regimes, but the total and bound PSDs do not. Depending on what particles the experimental measurement most represents, there are thus only a few restricted cases where one could see these two corners. Finally, depending on the ratio between the different time scales, the scaling laws probed may either be $1/f^{3/2}$ or $1/f^{2}$ at high frequencies.
IV Discussion and Conclusions
In this paper, we have derived expressions for the power spectrum of particle number fluctuations in a minimal confined system with adsorption and diffusion. By comparing our analytical results to simulated data, we have demonstrated that our expressions are valid over a wide range of parameters. We have investigated fluctuations in the number of bound, free, and total number of particles, finding remarkably different behavior across these quantities for a given parameter set. This highlights one of the main results of our investigation, namely, that the pure diffusion and pure adsorption components do not add up linearly in the PSD.
Spectra show a number of characteristic scalings that reflect key physical mechanisms and properties of the system, as outlined in Table 2. By considering the 1D case, we have also eliminated the effect of geometry on scaling Gravelle, Netz, and Bocquet (2019); Robin et al. (2023a) and significantly simplified interpretation of the results. Interestingly, our analytic results indicate that there are only a small number of robust scalings which emerge directly from physical mechanisms, with slopes of $-1/2$, $-3/2$, and $-2$ respectively. Observations of other scalings are the result of intermediate transitions between regimes or combination of two physical processes. The scaling of the power spectra thus also provides valuable insight into the dominant mechanism generating number fluctuations across different frequency ranges.
Scaling Range Mechanism
plateau low $f$ finite system size, finite global decorrelation time
$1/f^{1/2}$ intermediate $f$ diffusive re-entrance effects
$1/f^{3/2}$ high $f$ diffusive crossings between regions
$1/f^{2}$ high $f$ independent events, like adsorption/desorption
Table 2: Key scalings in the system with their corresponding timescales or range of frequencies and physical mechanism.
While our model is clearly simplistic, the diversity of results highlights how subtle and convoluted a PSD signal can be, and justifies the use of basic theoretical models to de-convolve PSD traces. Having at hand knowledge of when two corners and two slopes may be present, we can move on to investigate more complex systems. Such extensions could include more advanced geometries, which could be investigated for instance by adding another 1D line corresponding to the reservoirs Marbach (2021). Separately, we can carry out similar characterization of the 1D system after incorporation of other physical processes: for example, particle-particle interactions or a driving force in the channel. This would more accurately represent particles flowing through the channel under an applied electric field or pressure gradient, as is often the case in nanoporous or microfluidic experiments, and would open the door to probing non-equilibrium systems.
Real systems are clearly more complex than the situation considered here, but we can find scenarios in experiments that correspond to the different separation of timescales between diffusion and adsorption that we have studied. These are summarized in Fig. 7, left-hand side, and clearly span the entire phase diagram, indicating that there is no general shape to be expected from a PSD even with such simple ingredients. The ordering of slopes can also be used in the first instance to determine relative timescales of diffusion and adsorption: a relatively simple criterion to check. An interesting question for all of these cases is which of our modeled PSDs – total, free, or bound – best reflects the physical quantity measured experimentally. For example, nanopore systems are often investigated via ionic current fluctuations. Here, the current is a measure of the ions in solution traveling through the pore; therefore it could be reasonable to use the total PSD as an analogue for fluctuations in the current, assuming a degree of correlation between the pore conductance and total particles blocking the channel (adapter, sugar, etc) Secchi et al. (2016). In assays or nano-electrical devices, by design, the adsorbed particles may have the most significant effect on a macroscopic fluctuating signal, in which case the bound PSD is most relevant.
Our modeling does suggest a degree of caution in interpreting fluctuating behavior through only the knowledge of the PSD in experiments. In some cases, the difference between spectra are very small, e.g. between a pure diffusion spectrum and the total or free PSDs when all timescales are comparable, or between a pure adsorption spectrum and the free PSD when adsorption is much faster than desorption. Differences between separate scenarios are clear when we know the correct parameters of the system to fit and compare to the theoretical results. In real experimental data, this is likely to be much harder to do, especially when two or more timescales are comparable, and the situation may become an inverse problem where several parameter combinations produce similar power spectra. Nonetheless, we have identified that the two corners can only appear under a specific separation of timescales suggesting that when the two corners are seen, this means that one can in principle identify which regime of parameter space one is most likely in. This also suggests possible – although likely non trivial – experimental designs where e.g. tuning the diffusive timescale may bring the system in one or another regime.
While our focus has predominantly been on investigating the PSD, it is clear that it is not the only way, and sometimes not even the relevant way, to investigate a fluctuating signal. In particular, because the PSD is generally obtained by averaging the signal over rolling time windows, it “erases” any memory effects in the signal. Other calculations on the fluctuating signal have revealed the non-Markovian nature of processes Fuliński et al. (1998) but are only rarely used. Given the current enthusiasm for memristive nanofluidic devices, which are based on long-time memory effects Emmerich et al. (2024); Robin et al. (2023b); Kamsma et al. (2024), it is likely that more diverse investigations of noise signals will soon be necessary.
Acknowledgements.
The authors acknowledge fruitful discussions with Tristan Cerdin, Simon Gravelle and Pierre Levitz. The conception of this work was made possible thanks to numerous discussions with Aleksandar Donev. A. D. Y. is grateful to Max Howe for a very constructive discussion. S. M. received funding from the European Union’s Horizon 2020 research and innovation program under the Marie Skłodowska-Curie Grant Agreement No. 839225, MolecularControl. A. L. T. acknowledges funding from a Royal Society University Research Fellowship (URF/R1/211033) and from from EPSRC (EP/X02492X/1).
Author Declarations
Conflict of Interest
The authors have to conflicts to disclose.
Author Contributions
Anna Drummond Young: Conceptualization (equal); Investigation (lead); Methodology (equal); Software (lead); Visualization (equal); Writing - original draft (equal); Writing - review & editing (equal). Alice L. Thorneywork: Conceptualization (equal); Funding acquisition (lead); Methodology (equal); Supervision (equal); Visualization (equal); Writing - original draft (equal); Writing - review & editing (equal). Sophie Marbach: Conceptualization (equal); Investigation (equal); Methodology (equal); Supervision (equal); Visualization (equal); Writing - original draft (equal); Writing - review & editing (equal).
Data Availability Statement
The data that support the findings of this study are available within the article. To make the results in Eq. (12) more accessible, we have developed a sandbox where users can input their system parameters of choice and solve to find the resulting power spectra. The sandbox is in Google Colab, and available on GitHub: https://github.com/ady-100/Noise-in-Nanofluidics.
Appendix A Solving the rates problem
We designate $p_{C}$, $p_{L}$, and $p_{R}$ as representing variable $p$ in the channel (C) and left (L) and right (R) reservoirs, respectively. We define the Laplace transform $\hat{f}(x,s)$ of a function $f(x,t)$ as
[…]
In Laplace space, a time derivative is equivalent to multiplication by $s$, such that $\mathcal{L}\{\partial_{t}p\}(s)=s\hat{p}(s)-p(x,0)$. This notably leads to the incorporation of the initial conditions ($p(x,0)$ and $q(x,0)$) via the time derivative. Accordingly, we transform Eq. (1) into Laplace space, splitting $p$ up into $p_{C}$, $p_{L}$, and $p_{R}$ and using the initial conditions Eq. (2),
$\begin{cases}&\hat{p_{C}}=-\frac{k_{\rm on}}{s}\hat{p_{C}}+\frac{k_{\rm off}}{s}\hat{q}+\frac{D}{s}\partial_{xx}\hat{p_{C}}+\frac{p_{0}}{s}\\ &\hat{q}=\frac{k_{\rm on}}{s}\hat{p_{C}}-\frac{k_{\rm off}}{s}\hat{q}+\frac{q_{0}}{s}\\ &\hat{p_{L}}=\frac{D}{s}\partial_{xx}\hat{p_{L}}\\ &\hat{p_{R}}=\frac{D}{s}\partial_{xx}\hat{p_{R}}\end{cases}$ (20)
with boundary and continuity conditions
$\begin{cases}&\mathrm{at}\,\,x=-L/2\,\,\,\,\partial_{x}\hat{p_{L}}=0\\ &\mathrm{at}\,\,x=-L_{0}/2\,\,\,\,\displaystyle\begin{cases}\hat{p_{L}}=\hat{p_{C}}\\ \partial_{x}\hat{p_{L}}=\partial_{x}\hat{p_{C}}\end{cases}\\ &\mathrm{at}\,\,x=L_{0}/2\,\,\,\,\displaystyle\begin{cases}\hat{p_{R}}=\hat{p_{C}}\\ \partial_{x}\hat{p_{R}}=\partial_{x}\hat{p_{C}}\end{cases}\\ &\mathrm{at}\,\,x=L/2\,\,\,\,\partial_{x}\hat{p_{R}}=0.\end{cases}$ (21)
We solve the homogeneous equations for $\hat{p_{L}}$ and $\hat{p_{R}}$ by introducing the notation
$m_{0}(s)=\sqrt{\frac{s}{D}}$ (22)
to obtain
$\hat{p_{L}}$ $=A_{L}e^{m_{0}x}+B_{L}e^{-m_{0}x}$ (23)
$\hat{p_{R}}$ $=A_{R}e^{m_{0}x}+B_{R}e^{-m_{0}x}$
where $A_{L/R}$ and $B_{L/R}$ are constants to be determined from the boundary conditions. To solve for $\hat{p_{C}}$ and $\hat{q}$, we write $\hat{q}$ in terms of $\hat{p_{C}}$ using Eq. (20)
$\hat{q}=\frac{k_{\rm on}\hat{p_{C}}+q_{0}}{s+k_{\rm off}}$ (24)
which allows us to simplify the equation for $\hat{p_{C}}$ into
$\partial_{xx}\hat{p_{C}}-\frac{s(s+k_{\rm off}+k_{\rm on})}{D(s+k_{\rm off})}\hat{p_{C}}=-\frac{(k_{\rm off}(p_{0}+q_{0})+sp_{0})}{D(s+k_{\rm off})}.$ (25)
Introducing the notation
$m_{1}(s)=\sqrt{\frac{s(s+k_{\rm off}+k_{\rm on})}{D(s+k_{\rm off})}}$ (26)
we find the solution to Eq. (25)
$\hat{p_{C}}=A_{C}e^{m_{1}x}+B_{C}e^{-m_{1}x}+\frac{k_{\rm off}(p_{0}+q_{0})+sp_{0}}{s(s+k_{\rm off}+k_{\rm on})}$ (27)
where again $A_{C}$ and $B_{C}$ are constants to be determined from the boundary conditions.
Combining this with Eq. (24) we find:
$\hat{q}=\frac{k_{\rm on}}{s+k_{\rm off}}(A_{C}e^{m_{1}x}$ $+B_{C}e^{-m_{1}x})$ (28)
$+$ $\frac{k_{\rm on}(p_{0}+q_{0})+sq_{0}}{s(s+k_{\rm off}+k_{\rm on})}.$
There is a degree of symmetry between the results, and between the integration constants. Namely, using the boundary conditions in Eq. (21), we have
$A_{C}=B_{C}=-\frac{m_{0}(k_{\rm off}(p_{0}+q_{0})+p_{0}s)}{s(s+k_{\rm off}+k_{\rm on})}\times...$ (29)
$\frac{e^{L_{0}m_{1}/2}(e^{Lm_{0}}-e^{L_{0}m_{0}})}{(m_{0}-m_{1})(e^{Lm_{0}}-e^{L_{0}(m_{0}+m_{1})})+(m_{0}+m_{1})(e^{Lm_{0}+L_{0}m_{1}}-e^{L_{0}m_{0}})}.$
We do not report the expressions of the other integration constants, since they will not be required in the following steps.
Appendix B Further correlation functions
B.0.1 Number of adsorbed particles
Following the same method as Sec. II.2, we calculate the correlation function for the number of bound particles in the channel
$C_{\rm on}(t)$ $=\langle N_{\rm on}(t)N_{\rm on}(0)\rangle$ (30)
$=\mathcal{N}\psi_{\rm on\rightarrow on}(t)$
where $\psi_{\rm on\rightarrow on}(t)$ represents the probability that a particle starting adsorbed in the channel is still adsorbed in the channel at time $t$. Importantly, this means that the initial conditions in Eq. (2) now become $p_{0}=0$. We have
[…]
We take, as for the total number, the time derivative and Laplace transform and using Eqs. (27) and (28), we obtain:
$\hat{C}_{\rm on}$ $(s)=\frac{2\mathcal{N}}{s}\Bigg{[}\frac{sk_{\rm on}A_{C}}{m_{1}(s+k_{\rm off})}$ (32)
$\Big{(}e^{L_{0}m_{1}/2}-e^{-L_{0}m_{1}/2}\Big{)}-\frac{q_{0}k_{\rm off}L_{0}}{2(s+k_{\rm off}+k_{\rm on})}\Bigg{]}.$
B.0.2 Number of unadsorbed particles
Again we follow the same method as Sec. II.2. We have:
$C_{\rm off}(t)$ $=\langle N_{\rm off}(t)N_{\rm off}(0)\rangle$ (33)
$=\mathcal{N}\psi_{\rm off\rightarrow off}(t)$
where this time $\psi_{\rm off\rightarrow off}(t)$ represents the probability that a particle starting unadsorbed in the channel is still unadsorbed in the channel at time $t$. Importantly, this means that the initial conditions now include $q_{0}=0$. We have:
[…]
We take, as for the total number, the time derivative and Laplace transform and using Eqs. (27) and (28), we obtain:
$\hat{C}_{\rm off}$ $(s)=-2\mathcal{N}\Bigg{[}A_{C}\Big{(}e^{L_{0}m_{1}/2}-e^{-L_{0}m_{1}/2}\Big{)}$ (35)
$\Big{(}\frac{k_{\rm on}}{m_{1}(s+k_{\rm off})}-\frac{m_{1}}{m_{0}^{2}}\Big{)}+\frac{p_{0}k_{\rm on}L_{0}}{2s(s+k_{\rm off}+k_{\rm on})}\Bigg{]}.$
Appendix C Computational methods
The system we simulate is that described in Fig. 2. Particles are initialized along the lines $P$ and $Q$ with an initial distribution calculated from the expected number of particles in each region according to the probability densities $p_{0}$ and $q_{0}$. This allows the steady state to be reached faster than random initialization. The system contains $\mathcal{N}$ particles and $N_{\rm T}$ steps are carried out in total. We use $\mathcal{N}=100$ and $N_{\rm T}=2\times 10^{5}$. The time step used is $dt$, and is chosen to be sufficiently small compared to the timescales of diffusion, adsorption, and desorption that it captures their full behavior, i.e.:
$dt$ $\ll 1/k_{\rm on}$ (36)
$dt$ $\ll 1/k_{\rm off}$
$dt$ $\ll L_{0}^{2}/D.$
The probability for a given particle inside the channel to adsorb in a time step is $p_{\rm on}=k_{\rm on}dt$ and the probability to desorb is $p_{\rm off}=k_{\rm off}dt$.
Each step is carried out as follows. For each particle, we first check if it is already adsorbed. If it is, a random number is sampled from a uniform distribution $r_{1}\sim U(0,1)$. If $r_{1}<p_{\rm off}$, the particle unadsorbs, and if $r_{1}>p_{\rm off}$, it remains adsorbed. If the particle was not previously adsorbed, but is in the channel region, a second random number $r_{2}\sim U(0,1)$ is generated. If $r_{2}<p_{\rm on}$, the particle adsorbs, and if $r_{2}>p_{\rm on}$, it remains unadsorbed. If it remains unadsorbed, or if it was initially in the reservoirs, the particle moves according to:
$X(t+dt)=X(t)+\sqrt{2Ddt}\,\mathcal{W}_{t}$ (37)
where $\mathcal{W}_{t}$ is a sampled from a normal distribution at each time step, $\mathcal{W}_{t}\sim N(0,1)$. If $X(t+dt)$ lies inside the system it is accepted; however, if it lies outside the allowed region it is reflected back into the allowed region according to Snell’s law (specular reflection) Scala, Voigtmann, and De Michele (2007).
We measure the number of particles in the channel, or specific region of the channel, at every time step. To characterize these fluctuations, we calculate the PSD which can be defined as:
$S(f)=|\mathcal{F}\{x(t)\}|^{2}$ (38)
where $x(t)$ is the time-varying signal and $\mathcal{F}$ represents a Fourier transform. Since we are measuring a discrete signal – the particle number fluctuations – we cannot apply Eq. (38) directly. We instead estimate the power spectrum using Welch’s method Welch (2003), which takes the average PSD from multiple overlapping sections of the signal. This reduces statistical error on $S(f)$, an advantage over other methods of PSD estimation. To further reduce error, we run each simulation $10$ times and average the estimated PSDs. The spectrum is then smoothed by logarithmically binning along the frequency axis and averaging $S(f)$ in each bin. Then the PSD of the simulated number fluctuations can be compared directly to the theoretical results.
Appendix D Additional results in the full case
Here we consider additional representative results in the full case with adsorption processes and diffusion.
D.1 Diffusion Slower than Adsorption/Desorption
First we consider the complementary case to that in Sec. III.3.1. Changing the ordering of the adsorption and desorption timescales does not significantly affect the shape of the power spectra, but nonetheless there are some subtle changes. Results are in Fig. 8.
Figure 8: Slow diffusion, fast adsorption/desorption – Case 1 (continued): $\tau_{\rm cha}\gg\tau_{\rm ads}\sim\tau_{\rm des}$. Left: total PSD. Middle: bound PSD. Right: free PSD. We use $L=2\ell_{0}$, $k_{\rm on}=60/\tau_{0}$, and $k_{\rm off}=20/\tau_{0}$. Note that the main difference compared to Fig. 5 is that relative ordering of the rates of adsorption and desorption are now reversed, here $k_{\rm on}\gg k_{\rm off}$. Labels are similar to Fig. 5.
Now, when the adsorption rate is faster than desorption and binding is favored, the total PSD lies further from the purely diffusive line (dotted) at high frequencies. In this region the free PSD lies closer to the pure adsorption line (dash-dot). These changes can be reconciled with the fact that increasing the adsorption rate makes binding more favorable, so particles adsorb more quickly on entering the channel, on average. This suppresses the high frequency $-3/2$ scaling due to the diffusive motion of particles across boundaries in the total PSD, and also increases the importance of the adsorption process to the free PSD, causing its scaling to tend to $-2$.
Figure 9: All physical processes have comparable timescales – Case 3: $\tau_{\rm cha}\sim\tau_{\rm ads}\sim\tau_{\rm des}$. Left: total PSD. Middle: bound PSD. Right: free PSD. We use $L=2\ell_{0}$ and $k_{\rm on}=k_{\rm off}=0.4/\tau_{0}$.
D.2 Comparable Timescales
Spectra for the case where diffusion, adsorption, and desorption all have similar timescales is shown in Fig. 9. Here, the two corner frequencies and intermediate scaling we could observe clearly in the previous two cases collapse, with all three PSDs showing only one plateau and corner frequency. This illustrates how a complex physical situation with overlapping timescales could be difficult to interpret experimentally. While there are four timescales and two distinct processes (adsorption and diffusion) in the system, they are masked by the simple shape of the PSD.
However, our model and simulations provide a possible means of distinguishing this case from the similar-looking cases of pure diffusion and pure adsorption. The total PSD is close to the pure diffusive case but at intermediate frequencies transiently follows the adsorption case with a scaling of $-2$. The bound PSD, while similar to the case of pure adsorption, has a higher plateau. The free PSD resembles the case of pure diffusion, but its slope is steeper than $-3/2$ at high frequencies as it is still affected by adsorption.
Where could we observe this kind of behavior? One example is in the adsorption of polymers inside glass nanopores Knowles et al. (2021). We could also see an overlap in diffusion and adsorption in DNA hybridization microarray assays Gadgil et al. (2004). Although diffusion of small molecules is generally “fast”, the diffusive timescale becomes long when the channel length involved increases to hundreds of ${\rm nm}$ or even ${\rm\mu m}$.
Figure 10: Adsorption much faster than desorption – Case 4: $\tau_{\rm des}\gg\tau_{\rm sys}\sim\tau_{\rm cha}\gg\tau_{\rm ads}$. Left: total PSD. Middle: bound PSD. Right: free PSD. We use $L=2\ell_{0}$, $k_{\rm on}=4/\tau_{0}$, and $k_{\rm off}=0.04/\tau_{0}$.
D.3 Separation of Adsorption and Desorption Timescales
Lastly, we consider the case of the separation of adsorption and desorption timescales. When desorption is much faster than adsorption, we tend to the pure diffusive case, so we do not consider that here. However, when adsorption is the faster process, we have a more interesting case where particles enter the channel and immediately adsorb. This means particles effectively transition between the reservoirs and the bound state. Although this may seem like an extreme case, this situation can occur physically in any system where adsorption is heavily favored. This is the case in the catalysis of benzene alkylation in zeolites Zhao et al. (2016), as well as in microporous biomolecule sensors Hansen et al. (2009).
Representative results are shown in Fig. 10. The PSDs now have more complex shapes. The total PSD shows at several scaling regimes, with a corner at the diffusive timescale and a kink at the adsorption timescale. Above the adsorption timescale, it scales as approximately $-3/2$. The bound PSD also shows at least two separate scalings and corner frequencies. At high frequencies, it resembles the pure adsorption case with a $-2$ scaling. The free PSD is very close to the case of pure adsorption (telegraph noise) but its slope at high frequencies is less steep than $-2$. Interestingly, the diffusive timescale along the channel is the only timescale not corresponding to a change in slope for these PSDs - this may be because very few (if any) particles are able to diffuse along the channel without adsorption.